Giantage Corporation produces windows used in residential construction. The company expects to produce 44,550 units next year. With regards to direct materials, each unit of product requires 12 square feet of glass at a cost of $1.50 per square foot. With regards to direct labor, each unit of product requires 2 labor hours at a cost of $15/hour. The following information relates to the manufacturing overhead budget.

Variable Overhead Costs

Indirect materials

$2.50 per unit

Indirect labor

$3.20 per unit

Other

$1.70 per unit

Fixed Overhead Costs

Salaries

$50,000

Rent

$60,000

Depreciation

$36,000

  1. Calculate the direct materials budget for Giantage.
  2. Calculate the direct labor budget for Giantage.
  3. Calculate the manufacturing overhead budget for Giantage.
Answer #1

a. Direct material budget = 44550*12*1.50 = $801900

b. Direct labor budget = 44550*2*15 = $1336500

c. Manufacturing overhead budget = (2.50+3.20+1.70)*44550 + (50000+60000+36000)

= $475670